# Dark Matter (DM); Dark Stars (DSs); High-z galaxies; Supermassive Black Holes (SMBHs), Little Red Dots (LRDs); Blue Monsters; Over-massive BH galaxies (OBGs)

### v1

@@ 2.1 Dark Star Candidates Found
JWST is sensitive enough to detect Supermassive Dark Stars, even if unlensed (Ilie:2012), with the first ever Dark Star photometric candidates having been identified by (Ilie:2023JADES): JADES-GS-z11-0 , JADES-GS-z12-0 , and JADES-GS-z13-0 . In a followup study, ilie2025spectroscopicsupermassivedarkstar analyzed spectra and morphology for four of the most distant objects ever observed: JADES-GS-z11-0 , JADES-GS-z13-0 , JADES-GS-z14-0 , and JADES-GS-z14-1 and found that each of those objects are consistent with Supermassive Dark Stars powering a nebula; for a summary of values of best fit SMDSs model parameters for each of those objects see Table 1, where we use the short name for each object, i.e. we omit the JADES designation. SMDS mass ($M_{SMDS}$), estimated redshift ($z$), the density of H in the nebula surrounding the SMDSs ($n_{H}$), the effective size of the nebula ($r_{e}$), and the Sérsic Index are listed. As JADES-GS-z14-1 is consistent with a point object, it was modeled as a pure SMDS, by assuming negligible $n_{H}$. The best fit SMDSs SEDs are plotted against NIRSpec data in Fig. 1. More details regarding methodology of finding the best fit models, and explicitly showing how their morphology matches that inferred for those objects from NIRCam data can be found in ilie2025spectroscopicsupermassivedarkstar. For JADES-GS-z14-0 a tentative He II 1640 Å absorption feature, a smoking gun signature for dark stars, has been identified at a $S/N\sim 2$ level by ilie2025spectroscopicsupermassivedarkstar (see Fig. 2).
Table 1: Summary of best fit parameters for the four SMDSs spectroscopic candidates found in ilie2025spectroscopicsupermassivedarkstar.
@@
* For comparison, in the third column we list the estimated stellar mass ($M_{\star}$) if the same objects are modeled as regular galaxies, with no Dark Star component. (see JWSTBLueMonsters:2025A&A...694A.286F, Table 1)
@@
Figure 2: He II 1640 Å absorption feature identified in the spectrum of JADES-GS-z14-0 by ilie2025spectroscopicsupermassivedarkstar. Here we calculate the SNR of detection based on a polynomial fit (orange) to the observed spectrum (blue). The location of He II line is shown in black, and the size of the feature is shaded in gray. The feature is below the continuum beyond the level of noise. Estimated $SNR\simeq 2.31$.
[…]
Figure 3: He II 2511 Å absorption feature identified in the spectrum of JADES-GS-z13-0 . Estimated $SNR\simeq 4.5$.
@@ 3 Supermassive Dark stars at the core of the JWST’s Blue Monsters?
Despite their size ($\sim 10$ AU), SMDS can appear to be significantly more extended due to the effects of nebular emission (Zackrisson:2011; ilie2025spectroscopicsupermassivedarkstar). As such they can appear as marginally resolved objects in JWST, with effective sizes of up to $\sim 500$ parsecs (see Table 1). It is therefore possible that many of the “Blue Monster” objects found by JWST are in fact Dark Stars masquerading as galaxies. In fact, the four objects shown to be consistent with SMDSs in ilie2025spectroscopicsupermassivedarkstar, as discussed in Sec. 2, are a sub-sample of the 15 or so NIRSpec spectroscopically confirmed $z\gtrsim 10$ Blue Monsters (e.g. JWSTBLueMonsters:2025A&A...694A.286F, see Table 1 there). In a future publication we plan to analyze the data for the other 11, in terms of SMDSs or SDMSs embedded in proto-galaxies. Our expectation, based on the success rate of the SMDSs analysis for the objects we focused on so far, is that most of them would be consistent with SMDSs powering a nebula and/or SMDSs embedded in early proto-galaxies. Below we discuss the challenges posed by the the $z\gtrsim 10$ JWST ultra-compact, yet very bright, JWST Blue Monsters to standard models of formation of the first stars and galaxies in the Universe, and then explain how each of those problems is in fact a feature, once one adopts a Supermassive Dark Star interpretation of those objects.
@@
First, Blue Monsters are very compact, and based on their brightness, if interpreted as galaxies, they would require $\gtrsim 10^{8}M_{\odot}$ stars to be packed in a few hundred parsecs (see Table. 1 in JWSTBLueMonsters:2025A&A...694A.286F, or for the case of JADES-GS-z11-0 , JADES-GS-z13-0 , JADES-GS-z14-0 , and JADES-GS-z14-1 Table 1 here). As galaxies, this would require an extremely dense packing of stars within a few hundred parsecs. Some of those objects are not even resolved with JWST, implying sizes $\lesssim 100$ pc, further exacerbating the problem explained above. However, if interpreted as isolated Supermassive Dark Stars, potentially powering a nebula (as done in ilie2025spectroscopicsupermassivedarkstar) compactness is a feature of the model.
@@
Second, spectroscopically confirmed $z\gtrsim 10$ Blue Monsters show very low dust attenuation and dust-to-stellar mass ratios ($\xi_{d}$) inconsistent (by at least two orders of magnitude) with models that include standard dust production and no outflows, dominated by supernova dust production over destruction. Thus, their transparency is hard to reconcile with dust physics alone without invoking additional mechanisms (JWSTBLueMonsters:2025A&A...694A.286F). The standard solution to this problem is the Attenuation Free Model (AFM), where dust is produced in standard amounts but is pushed to kiloparsec scales by radiation-driven outflows during super-Eddington phases (e.g. BlueMonsters:2023MNRAS.520.2445Z). Conversely, (JWSTBLueMonsters:2025A&A...694A.286F) tested a “dust-free” alternative by exploring whether reduced net SN dust yield and/or enhanced interstellar destruction can lower dust enough to match observations. Even allowing wide parameter variations, a no-outflow model cannot reproduce the very low dust-to-stellar mass ratios implied by JWST spectra. However, SMDSs are perfect inhibitors of dust production. As they convert large amounts of gas into a single supermassive star which eventually collapses to a SMBH, i.e. leaving no dust behind. Therefore, when compared to the standard AFM solution, a Dark Star interpretation of the Blue Monsters does not require super-Eddington phases capable of pushing dust to kilo-parsec scales. Again, what is observed and apparently a puzzle, is just a feature of Supermassive Dark Stars.
@@
Lastly, if interpreted as galaxies, those Blue Monsters have to have converted gas to stars at rates sometimes exceeding 50% (e.g. Boylan-Kolchin:2023). This is in stark contrast with expectations based on numerical simulations of the formation of the first stars and galaxies, and it exceeds the conversion rate observed in the “local” universe by a factor of five or more. This problem can be quantified in terms of the specific star formation rate (SSFR), which, for the spectroscopically confirmed JWST galaxies at $z\gtrsim 10$ is remarkably high (e.g. BlueMonsters:2023MNRAS.520.2445Z; BlueMonstersNumbers:2024ApJ...965...98C; JWSTBLueMonsters:2025A&A...694A.286F). On the other hand, a SMDSs interpretation can alleviate this problem in two ways. First, SMDSs shine almost at the Eddington limit, i.e. attaining the lowest theoretical limit for the mass to light ratio of a a system. This is evidenced, for example in Table 1, where the stellar mass ($M_{\star}$) required in order to explain the four Blue Monsters discussed there as galaxies is two-tree orders of magnitude higher than the mass of the corresponding SMDSs model ($M_{SMDS}$) for each object. Therefore, within a SMDSs interpretation, the inferred star formation rates would be much lower than those for obtained from the same $z\gtrsim 10$ JWST data, under a galaxy interpretation. Moreover, SMDSs are cool (i.e. $T_{eff}\lesssim 50,000~K$), and, as such, they can accrete the baryons surrounding them almost indefinitely (Freese:2010smds). Whenever a SMDSs runs out of DM fuel, or if it grows sufficiently massive, and therefore triggers general relativistic instabilities, it will undergo a core collapse.  [^4] In Freese2025_DarkStarInstability we find that the collapse is triggered, depending on the mass of the WIMP powering a SMDS, somewhere in the $M_{SMDS}\sim 10^{5}-10^{6}M_{\odot}$ range. As we shall see in the next section, this provides a natural mechanism for seeding SMBHs.
@@ 4 Two kinds of “heavy seeds:" Supermassive Dark Stars and DCBHs
Direct Collapse Black Holes (DCBHs) (e.g. Loeb:1994wv; Belgman:2006; Lodato:2006hw; Natarajan:2017; barrow:2018; Whalen:2020; Inayoshi:2020) are a well known mechanism of seeding OBHs. This theoretical scenario relies on nuclear burning stars entering a supergiant cool phase whenever the accretion rate exceeds $0.1~M_{\odot}\mathrm{yr}^{-1}$ (e.g. Hosokawa_2012). Thus, a very massive nuclear burning star can be born. One of the drawbacks of this mechanism is how unlikely the conditions to form a DCBH are (Natarajan:2017; DCBHsDensity), as they require the need for a very efficient companion star forming Dark Mater halo in close proximity in order to destroy molecular hydrogen clouds inside the DCBH host halo (Belgman:2006; Kiyuna_2023). In contrast, SMDS are able to form over a wider variety of redshifts, and a sizable fraction of DM halos would be able to seed SMDS (e.g. Ilie:2012; Banik:2019). For more details see Appendix C.
Following  Bogdan:2023UHZ1 we will focus on UHZ1, a JWST galaxy at $z\sim 10$, harboring a very bright quasar ($L_{bolo}\sim 5\times 10^{45}~\mathrm{erg\,s^{-1}}$). UHZ1 is the most striking example of the puzzle described above, in view of its extreme distance ($z\sim 10$) and inferred black hole mass ($M_{BH}\sim 10^{7}M_{\odot}$). Bogdan:2023UHZ1; Natarajan:2023UHZ1 conclude that UHZ1 is the best evidence so far for the DCBH scenario. Here, we propose an alternative interpretation, by showing that the SMBH powering the X-ray observed spectrum of UHZ1 could be the BH remnant of a Supermassive Dark Star.
Figure 4: BHs with masses between $10^{4}$ and $10^{5}M_{\odot}$, generated at $z\simeq 25$ and growing at the Eddington rate (tan shaded band), can explain the mass of UHZ1 (Solid black line) and the three previously known highest redshift quasars (denoted by diamond symbols at $z\sim 7.5$). A radiative efficiency of the accretion process $\eta=0.114$ is needed to reproduce the growth curves from Fig. 4 of Bogdan:2023UHZ1 (shaded tan band in this figure).
Below we summarize the main evidence for the “heavy seed” hypothesis as an explanation for UHZ1 (Bogdan:2023UHZ1; Natarajan:2023UHZ1). In Fig. 4 we plot, following Bogdan:2023UHZ1, the growth curves (tan band) for hypothetical supermassive black holes that have been seeded at $z\simeq 25$ with masses in the expected DCBH initial mass range: $10^{4}-10^{5}M_{\odot}$. As the tan band in Fig. 4 passes through the inferred BH mass for UHZ1, the authors of Natarajan:2023UHZ1 claim that this is the best evidence so far for the existence of DCBHs. Note that SMDSs could collapse to BHs of the same mass range at the same redshift. Therefore, we argue that in fact UHZ1 is in fact the best evidence so far for the need of “heavy seed” Black Holes at high redshifts, rather than for the DCBH scenario itself. In order to generate the growth curves (i.e. the orange band in the figure) we solve:
@@
Typical expected values for this parameter are $\eta\sim 0.1$, with the fiducial value of $0.057$ corresponding to accretion onto non-rotating BHs (Bardeen:1970BHs), while for maximally rotating BHs $\eta=0.43$ (Bardeen:1970BHs). Note that if we were to use $\eta=0.057$ in Fig. 4, the curves would have overshot the mass of UHZ1 and the $z\simeq 7.5$ quasars by orders of magnitude. We find that for $\eta=0.114$ we are able to reproduce the growth rates of Fig. 4 from Bogdan:2023UHZ1. Note that even a small change in $\eta$ would imply a significant change in the estimated BH mass at $z\sim 10$ or below. Therefore, quite a large degree of fine-tuning is necessary to explain UHZ1 and the other three $z\simeq 7.5$ quasars  [^5] The mass estimates of J0313-1806 (J0313), J1342+0928 (J1342+0928), and J1007+2115 (Yang2020) are much more accurate than that of UHZ1 (Bogdan:2023UHZ1), for which no BLR emission is detected. in terms of DCBHs formed at $z\simeq 25$. However, the degeneracy between $\eta$ and the DCBH formation redshift alleviates this problem.
As discussed in Sec. 2, Dark Stars are expected to form at the centers of DM minihalos at redshifts $z_{form}\in[30,10]$. They are extremely puffy, with radii of the order of a few A.U. Their initial masses are $\sim 1M_{\odot}$ and they grow via accretion for as long as they are powered by DM annihilation and for as long as there is material to accrete (Freese:2010smds). The accretion rate can be modeled with (constant) values typically ranging between $10^{-3}$ and $10^{-1}M_{\odot}\mathrm{yr}^{-1}$. In previous studies of dark stars (Spolyar:2009; Freese:2010smds) the physically motivated variable accretion rates of TanMcKee:2004; mckee2008formation were used as well, confirming results based on constant accretion rates, as long as the Dark Star remains cooler than $T_{eff}\sim 50,000$ K, which is the case for all the SMDSs models studied so far in the literature. It is remarkable that the evolution along the HR diagram, and therefore the observable properties of Dark Stars powered by Adiabatically Contracted DM are insensitive to the accretion rate. Note that even for the highest constant accretion rate considered here ($10^{-1}M_{\odot}\mathrm{yr}^{-1}$), the luminosity due to accretion onto a DS remains sub-Eddington, due primarily to the large radii of those objects.
@@
JWST IR data suggests that UHZ1 harbors a significant stellar population, with mass $\sim 10^{7}M_{\odot}$. This is actually a feature of the DCBH interpretation for UHZ1 (Bogdan:2023UHZ1; Natarajan:2023UHZ1), since a large Pop III stellar population is needed as a catalyst for the formation of the DCBH in a nearby satellite halo (for more details see Appendix C). In turn, there are a variety of ways to explain the UHZ1 data using Dark Stars, the simplest of which is a SMBH SMDS system. However, in view of the extended morphology of the UHZ1 system this simple interpretation is unlikely. The second possible explanation of the data is in fact very similar to the one proposed by Refs. Bogdan:2023UHZ1; Natarajan:2009. Namely, the system is a galaxy that harbors a SMBH of $M_{SMBH}\sim 10^{7}M_{\odot}$. In our scenario the seed of the Supermassive Black Hole is a Dark Star rather than a DCBH.

### v2

@@ 2.1 Dark Star Candidates Found
Table 1: Summary of best fit parameters for the four SMDSs spectroscopic candidates found in Ilie et al. (2025). Note that we assumed, for simplicity, that Dark Stars are powered by 100 GeV WIMPs, annihilating at the canonical $\langle\sigma v\rangle=3\times 10^{-26}cm^{3}/s$. In future work we plan to include the mass of the DM particle as a free parameter, which could be constrained based on observable properties of SMDSs.
@@
* For comparison, in the third column we list the estimated stellar mass ($M_{\star}$) if the same objects are modeled as regular galaxies, with no Dark Star component. (see Ferrara et al., 2025, Table 1)
JWST is sensitive enough to detect Supermassive Dark Stars, even if unlensed (Ilie et al., 2012), with the first ever Dark Star photometric candidates having been identified by (Ilie et al., 2023): JADES-GS-z11-0 , JADES-GS-z12-0 , and JADES-GS-z13-0 . In a followup study, Ilie et al. (2025) analyzed spectra and morphology for four of the most distant objects ever observed: JADES-GS-z11-0 , JADES-GS-z13-0 , JADES-GS-z14-0 , and JADES-GS-z14-1 and found that each of those objects are consistent with Supermassive Dark Stars powering a nebula; for a summary of values of best fit SMDSs model parameters for each of those objects see Table 1, where we use the short name for each object, i.e. we omit the JADES designation. SMDS mass ($M_{SMDS}$), estimated redshift ($z$), the density of H in the nebula surrounding the SMDSs ($n_{H}$), the effective size of the nebula ($r_{e}$), and the Sérsic Index are listed. As JADES-GS-z14-1 is consistent with a point object, it was modeled as a pure SMDS, by assuming negligible $n_{H}$. The best fit SMDSs SEDs are plotted against NIRSpec data in Fig. 1333For galactic fits see Hainline et al. (2024) (Figs. 3 and 4 there) presenting fits for JADES-GS-z11-0 and JADES-GS-z13-0 ) and Carniani et al. (2024) (extended data Figs. 9 and 10) for fits of JADES-GS-z14-1 and JADES-GS-z14-0 . At the level of the continuum the available NIRSpec data cannot be used to disambiguate between a galactic and a SMDSs interpretation for those four objects. More details regarding methodology of finding the best fit models, and explicitly showing how their morphology matches that inferred for those objects from NIRCam data can be found in Ilie et al. (2025). For JADES-GS-z14-0 a tentative He II 1640 Å absorption feature, a smoking gun signature for dark stars, has been identified at a $S/N\sim 2$ level by Ilie et al. (2025) (see Fig. 2).
[…]
@@
Figure 2: He II 1640 Å absorption feature identified in the spectrum of JADES-GS-z14-0 by Ilie et al. (2025). Here we calculate the SNR of detection based on a polynomial fit (orange) to the observed spectrum (blue). Namely, $SNR=D_{feature}/\sigma_{cont}$, where $D_{feature}$ represents the depth of the feature below the modeled continuum, which has a scatter (noise) quantified by $\sigma$. The location of He II line is shown in black, and the size of the feature is shaded in gray. The feature is below the continuum beyond the level of noise. Estimated $SNR\simeq 2.31$.
Figure 3: Absorption features identified in the spectrum of JADES-GS-z13-0 . Left panel: He II 1440 Å Estimated $SNR\simeq 2.2$. Right panel: He II 2511 Å. Esimated $NSR\simeq 4.6$.
@@ 3 Supermassive Dark stars at the core of the JWST’s Blue Monsters?
Despite their size ($\sim 10$ AU), SMDS can appear to be significantly more extended due to the effects of nebular emission (Zackrisson, 2011; Ilie et al., 2025). As such they can appear as marginally resolved objects in JWST, with effective sizes of up to $\sim 500$ parsecs (see Table 1). It is therefore possible that many of the “Blue Monster” objects found by JWST are in fact Dark Stars masquerading as galaxies. In fact, the four objects shown to be consistent with SMDSs in Ilie et al. (2025), as discussed in Sec. 2, are a sub-sample of the 15 or so NIRSpec spectroscopically confirmed $z\gtrsim 10$ Blue Monsters (e.g. Ferrara et al., 2025, see Table 1 there). In a future publication we plan to analyze the data for the other 11, in terms of SMDSs or SDMSs embedded in proto-galaxies. Our expectation, based on the success rate of the SMDSs analysis for the objects we focused on so far, is that most of them would be consistent with SMDSs powering a nebula and/or SMDSs embedded in early proto-galaxies. Below we discuss the challenges posed by the the $z\gtrsim 10$ JWST ultra-compact, yet very bright, JWST Blue Monsters to standard models of formation of the first stars and galaxies in the Universe, and then explain how each of those problems is in fact a feature, once one adopts a Supermassive Dark Star interpretation of those objects.
@@
First, Blue Monsters are very compact, and based on their brightness, if interpreted as galaxies, they would require $\gtrsim 10^{8}M_{\odot}$ stars to be packed in a few hundred parsecs (see Table. 1 in Ferrara et al. (2025), or for the case of JADES-GS-z11-0 , JADES-GS-z13-0 , JADES-GS-z14-0 , and JADES-GS-z14-1 Table 1 here). As galaxies, this would require an extremely dense packing of stars within a few hundred parsecs. Some of those objects are not even resolved with JWST, implying sizes $\lesssim 100$ pc, further exacerbating the problem explained above. However, if interpreted as isolated Supermassive Dark Stars, potentially powering a nebula (as done in Ilie et al. (2025)) compactness is a feature of the model.
@@
Second, spectroscopically confirmed $z\gtrsim 10$ Blue Monsters show very low dust attenuation and dust-to-stellar mass ratios ($\xi_{d}$) inconsistent (by at least two orders of magnitude) with models that include standard dust production and no outflows, dominated by supernova dust production over destruction. Thus, their transparency is hard to reconcile with dust physics alone without invoking additional mechanisms (Ferrara et al., 2025). The standard solution to this problem is the Attenuation Free Model (AFM), where dust is produced in standard amounts but is pushed to kiloparsec scales by radiation-driven outflows during super-Eddington phases (e.g. Ziparo et al., 2023). Conversely, (Ferrara et al., 2025) tested a “dust-free” alternative by exploring whether reduced net SN dust yield and/or enhanced interstellar destruction can lower dust enough to match observations. Even allowing wide parameter variations, a no-outflow model cannot reproduce the very low dust-to-stellar mass ratios implied by JWST spectra. However, SMDSs are perfect inhibitors of dust production. As they convert large amounts of gas into a single supermassive star which eventually collapses to a SMBH, i.e. leaving no dust behind. Therefore, when compared to the standard AFM solution, a Dark Star interpretation of the Blue Monsters does not require super-Eddington phases capable of pushing dust to kilo-parsec scales. Again, what is observed and apparently a puzzle, is just a feature of Supermassive Dark Stars.
@@
Lastly, if interpreted as galaxies, those Blue Monsters have to have converted gas to stars at rates sometimes exceeding 50% (e.g. Boylan-Kolchin, 2023). This is in stark contrast with expectations based on numerical simulations of the formation of the first stars and galaxies, and it exceeds the conversion rate observed in the “local” universe by a factor of five or more. This problem can be quantified in terms of the specific star formation rate (SSFR), which, for the spectroscopically confirmed JWST galaxies at $z\gtrsim 10$ is remarkably high (e.g. Ziparo et al., 2023; Casey et al., 2024; Ferrara et al., 2025). On the other hand, a SMDSs interpretation can alleviate this problem in two ways. First, SMDSs shine almost at the Eddington limit, i.e. attaining the lowest theoretical limit for the mass to light ratio of a a system. This is evidenced, for example in Table 1, where the stellar mass ($M_{\star}$) required in order to explain the four Blue Monsters discussed there as galaxies is two-three orders of magnitude higher than the mass of the corresponding SMDSs model ($M_{SMDS}$) for each object. Therefore, within a SMDSs interpretation, the inferred star formation rates would be much lower than those for obtained from the same $z\gtrsim 10$ JWST data, under a galaxy interpretation. Moreover, SMDSs are cool (i.e. $T_{eff}\lesssim 50,000~K$), and, as such, they can accrete the baryons surrounding them almost indefinitely (Freese et al., 2010). Whenever a SMDSs runs out of DM fuel, or if it grows sufficiently massive, and therefore triggers general relativistic instabilities, it will undergo a core collapse.  [^5] In Freese et al. (2025) we find that the collapse is triggered, depending on the mass of the WIMP powering a SMDS, somewhere in the $M_{SMDS}\sim 10^{5}-10^{6}M_{\odot}$ range. As we shall see in the next section, this provides a natural mechanism for seeding SMBHs.
@@ 4 Two kinds of “heavy seeds:" Supermassive Dark Stars and DCBHs
Direct Collapse Black Holes (DCBHs) (e.g. Loeb and Rasio, 1994; Begelman et al., 2006; Lodato and Natarajan, 2006; Natarajan et al., 2017; Barrow et al., 2018; Whalen et al., 2020; Inayoshi et al., 2020) are a well known mechanism of seeding OBHs. This theoretical scenario relies on nuclear burning stars entering a supergiant cool phase whenever the accretion rate exceeds $0.1~M_{\odot}\mathrm{yr}^{-1}$ (e.g. Hosokawa et al., 2012). Thus, a very massive nuclear burning star can be born. One of the drawbacks of this mechanism is how unlikely the conditions to form a DCBH are (Natarajan et al., 2017; Habouzit et al., 2016), as they require the need for a very efficient companion star forming Dark Mater halo in close proximity in order to destroy molecular hydrogen clouds inside the DCBH host halo (Begelman et al., 2006; Kiyuna et al., 2023). In contrast, SMDS are able to form over a wider variety of redshifts, and a sizable fraction of DM halos would be able to seed SMDS (e.g. Ilie et al., 2012; Banik et al., 2019). For more details see Appendix C.
Following  Bogdán et al. (2024) we will focus on UHZ1, a JWST galaxy at $z\sim 10$, harboring a very bright quasar ($L_{bolo}\sim 5\times 10^{45}~\mathrm{erg\,s^{-1}}$). UHZ1 is the most striking example of the puzzle described above, in view of its extreme distance ($z\sim 10$) and inferred black hole mass ($M_{BH}\sim 10^{7}M_{\odot}$). Bogdán et al. (2024); Natarajan et al. (2024) conclude that UHZ1 is the best evidence so far for the DCBH scenario. Here, we propose an alternative interpretation, by showing that the SMBH powering the X-ray observed spectrum of UHZ1 could be the BH remnant of a Supermassive Dark Star.
Figure 4: BHs with masses between $10^{4}$ and $10^{5}M_{\odot}$, generated at $z\simeq 25$ and growing at the Eddington rate (tan shaded band), can explain the mass of UHZ1 (Solid black line) and the three previously known highest redshift quasars (denoted by diamond symbols at $z\sim 7.5$). A radiative efficiency of the accretion process $\eta=0.114$ is needed to reproduce the growth curves from Fig. 4 of Bogdán et al. (2024) (shaded tan band in this figure).
Below we summarize the main evidence for the “heavy seed” hypothesis as an explanation for UHZ1 (Bogdán et al., 2024; Natarajan et al., 2024). In Fig. 4 we plot, following Bogdán et al. (2024), the growth curves (tan band) for hypothetical supermassive black holes that have been seeded at $z\simeq 25$ with masses in the expected DCBH initial mass range: $10^{4}-10^{5}M_{\odot}$. As the tan band in Fig. 4 passes through the inferred BH mass for UHZ1, the authors of Natarajan et al. (2024) claim that this is the best evidence so far for the existence of DCBHs. Note that SMDSs could collapse to BHs of the same mass range at the same redshift. Therefore, we argue that in fact UHZ1 is in fact the best evidence so far for the need of “heavy seed” Black Holes at high redshifts, rather than for the DCBH scenario itself. In order to generate the growth curves (i.e. the orange band in the figure) we solve:
@@
Typical expected values for this parameter are $\eta\sim 0.1$, with the fiducial value of $0.057$ corresponding to accretion onto non-rotating BHs (Bardeen, 1970), while for maximally rotating BHs $\eta=0.43$ (Bardeen, 1970). Note that if we were to use $\eta=0.057$ in Fig. 4, the curves would have overshot the mass of UHZ1 and the $z\simeq 7.5$ quasars by orders of magnitude. We find that for $\eta=0.114$ we are able to reproduce the growth rates of Fig. 4 from Bogdán et al. (2024). Note that even a small change in $\eta$ would imply a significant change in the estimated BH mass at $z\sim 10$ or below. Therefore, quite a large degree of fine-tuning is necessary to explain UHZ1 and the other three $z\simeq 7.5$ quasars  [^6] The mass estimates of J0313-1806 (Wang et al., 2021), J1342+0928 (Bañados et al., 2017), and J1007+2115 (Yang et al., 2020) are much more accurate than that of UHZ1 (Bogdán et al., 2024), for which no BLR emission is detected. in terms of DCBHs formed at $z\simeq 25$. However, the degeneracy between $\eta$ and the DCBH formation redshift alleviates this problem.
As discussed in Sec. 2, Dark Stars are expected to form at the centers of DM minihalos at redshifts $z_{form}\in[30,10]$. They are extremely puffy, with radii of the order of a few A.U. Their initial masses are $\sim 1M_{\odot}$ and they grow via accretion for as long as they are powered by DM annihilation and for as long as there is material to accrete (Freese et al., 2010). The accretion rate can be modeled with (constant) values typically ranging between $10^{-3}$ and $10^{-1}M_{\odot}\mathrm{yr}^{-1}$. In previous studies of dark stars (Spolyar et al., 2009; Freese et al., 2010) the physically motivated variable accretion rates of Tan and McKee (2004); McKee and Tan (2008) were used as well, confirming results based on constant accretion rates, as long as the Dark Star remains cooler than $T_{eff}\sim 50,000$ K, which is the case for all the SMDSs models studied so far in the literature. It is remarkable that the evolution along the HR diagram, and therefore the observable properties of Dark Stars powered by Adiabatically Contracted DM are insensitive to the accretion rate. Note that even for the highest constant accretion rate considered here ($10^{-1}M_{\odot}\mathrm{yr}^{-1}$), the luminosity due to accretion onto a DS remains sub-Eddington, due primarily to the large radii of those objects.
@@
JWST IR data suggests that UHZ1 harbors a significant stellar population, with mass $\sim 10^{7}M_{\odot}$. This is actually a feature of the DCBH interpretation for UHZ1 (Bogdán et al., 2024; Natarajan et al., 2024), since a large Pop III stellar population is needed as a catalyst for the formation of the DCBH in a nearby satellite halo (for more details see Appendix C). In turn, there are a variety of ways to explain the UHZ1 data using Dark Stars, the simplest of which is a SMBH SMDS system. However, in view of the extended morphology of the UHZ1 system this simple interpretation is unlikely. The second possible explanation of the data is in fact very similar to the one proposed by Refs. Bogdán et al. (2024); Natarajan et al. (2009). Namely, the system is a galaxy that harbors a SMBH of $M_{SMBH}\sim 10^{7}M_{\odot}$. In our scenario the seed of the Supermassive Black Hole is a Dark Star rather than a DCBH.